# Efficient Offline Reinforcement Learning: The Critic is Critical

### v1

@@ 6.3 Results and Analysis
door-human $0.6\pm 0.7$ $6.6\pm 6.4$ $0.1\pm 0.1$ $-0.2\pm 0.1$ $0.6\pm 0.6$ $\mathbf{13.2\pm 3.8}$
@@
door-cloned $0.3\pm 0.5$ $\mathbf{6.6\pm 3.6}$ $0.1\pm 0.5$ $-0.1\pm 0.0$ $0.0\pm 0.3$ $0.2\pm 0.5$
@@
Table 2: Normalized average returns ($\pm$ standard deviation) on Adroit, by averaging performance between 200k and 300k updates over 4 independent seeds. Our combined algorithms ensure that both the actor and critic are regularized to stay close to the behavior policy after pre-training, often giving greater performance than the component algorithms when learning from limited demonstrations.
@@
We find that behavior cloning (BC) provides a strong baseline, notably greatly outperforming previously quoted BC performances on these environments due to our addition of LayerNorm. We see that CQL and EDAC (with LayerNorm) perform reasonably with this evaluation approach using the same hyperparameters as specified in the original papers (Kumar et al.,, 2020; An et al.,, 2021), although may benefit from a greater training budget. The original TD3+BC paper did not consider the Adroit environments, but we see comparable performances with relatively strong behavior cloning regularization (tuned with $\alpha=1$ for the pen environments and $\alpha=0.1$ for other environments). However, our additions of CQL regularization to TD3+BC and BC regularization to EDAC both generally lead to improved performance in these environments. In fact, we find that our new hybrid algorithms perform similarly using this evaluation procedure even without pre-training, due to their strong inherent regularization, as shown and discussed further in Appendix I. Performance plots in these environments for the human datasets are provided in Appendix J.
7 Conclusion
We have demonstrated that pre-training policies and value functions to first be consistent with the provided offline dataset can improve the efficiency and stability of subsequent off-policy reinforcement learning. In particular, this consistency can reduce subsequent inefficiency and instability associated with bootstrapped temporal difference learning, and can more than halve the number of updates (and therefore computation time) required for state-of-the-art offline algorithms to converge on standard environments. Building on our insight into the importance of policy and value consistency, we demonstrated that combining regularization on both components can enable greater performance than policy or value regularization alone when learning from limited human demonstrations. We hope our research inspires further work towards bridging the gap between classic and deep algorithms to improve the efficiency and stability of offline reinforcement learning as scale continues to increase.
Acknowledgments and Disclosure of Funding
Appendix A Additional Related Work on Use of Monte-Carlo Values and Pre-Training in Reinforcement Learning
Recent work (Fujimoto et al.,, 2022; Patterson et al.,, 2022; Chen and Jiang,, 2019) has demonstrated that the Bellman error can be a poor proxy for the real value error, particularly when used for incomplete, off-policy datasets as in the offline setting, causing significant issues with utilising the Bellman error as the objective for training value functions offline. Monte-Carlo (MC) return estimates have previously been used successfully in online reinforcement learning to improve the sample efficiency of online exploration (Bellemare et al.,, 2016; He et al.,, 2016; Ostrovski et al.,, 2017; Oh et al.,, 2018; Tang,, 2021; Wilcox et al.,, 2022), but none of these approaches consider how to use MC returns in offline reinforcement learning, which is becoming of increasing importance for scaling reinforcement learning (Kumar et al., 2022a, ).
In the context of offline RL, Brandfonbrener et al., (2021) similarly recognise the effectiveness of learning the behaviour policy and value function before improvement, but do not consider the use of supervised learning to improve efficiency, and only take one step of TD improvement to prevent out-of-distribution extrapolation of this value function, rather than the more general and controllable combination of actor and critic regularization we propose. Pre-training policies with imitation learning for offline RL was recently investigated by Orsini et al., (2021), but they found that the gain from pre-training is generally insignificant due to policy updates from randomly initialised critic networks causing the policy to rapidly deteriorate (as we find in Section 5), motivating our work on pre-training the critic. Other work has considered pre-training off-policy algorithms using expert demonstrations (Goecks et al.,, 2020; Zhang and Ma,, 2018), but only consider the online setting and do not consider efficiency. Our work provides the first analysis of the benefits of pre-training with a supervised value-error objective, leading to more efficient and stable subsequent off-policy reinforcement learning.
Appendix B Rational for Separation of Actor and Critic Pre-training for Entropy regularized Reinforcement Learning
In section 3.3 and algorithm 2 of the main text, we propose separating the pre-training of the policy and value network into two separate phases for entropy-regularized reinforcement learning algorithms. By first pre-training the policy with soft behaviour cloning, an approximate behaviour policy can be learned which then enables approximate behaviour entropy bonuses to be included in the subsequent pre-training of the critic. However, an alternative approach could be to pre-train the policy and critic in parallel, as in the pure return maximisation framework. In theory this would require updating the returns-to-go for each policy update to incorporate the changing entropy bonuses as the variance of the policy is updated. Since this requires a complete forward pass of the policy for all subsequent states in the trajectories of the states sampled for an update, this makes the pre-training infeasibly expensive.
Another potential approach is to only train the mean of the standard Gaussian policy to match the behaviour policy, and keep the variance constant such that all entropy bonuses could be caluclated an intialisation and would be unaffected by policy pre-training. However, we note that the standard tanh squashing applied to the policy to keep the sampled action within the environment action bounds leads to a changing entropy of the resulting policy, even with the Gaussian variance kept constant.
A final approach we considered was to compute the soft returns-to-go based on the initialisation policy, and then only pre-train the values (no behaviour cloning). While this approach was successful and led to training efficiency gains, the rapid updating of the actor at the beginning of training (and particularly the rapidly changing policy entropy) quickly leads to inconsistent values, so we found that the investment in pre-training the policy with soft behaviour cloning first was worth the computational time in most cases.
Appendix C Investigation Into Affect of LayerNorm
Figure 3: Investigation into the affect of adding LayerNorm to both the actor and critic networks for TD3+BC on Hopper and HalfCheetah-medium. All lines show mean and standard deviation in normalised return at each timestep over 3 seeds.
We investigate the effect of the addition of LayerNorm (Ba et al.,, 2016) to both the actor and critic networks for TD3+BC on the Hopper-medium and HalfCheetah-medium datasets. The standard author implementation of TD3+BC (along with that of SAC-N, EDAC and most other off-policy reinforcement learning algorithms (Tarasov et al.,, 2022)) does not include any form of representational normalisation, and is shown in green. We consider the addition of LayerNorm after every linear layer in the network (before activation) except the final linear layer. We find that adding LayerNorms to the critic network leads to significant improvement in training efficiency and stability. This independently verifies the findings of Ball et al., (2023), who hypothesise that this occurs because the normalisation prevents severe value extrapolation for out-of-distribution actions, leading to overestimation error. Surprisingly, we find that the addition of LayerNorm to the actor (without addition to the critic) leads to worse efficiency and stability than the default (no LayerNorm) for the HalfCheetah environment. However, the addition of LayerNorm to both the actor and the critic leads to greater training efficiency and stability than the default or applying either normalisation alone, for both environments.
We find that these insights generally hold across investigated environments and datasets. However, we notice this this addition comes at the cost of a few percent in final performance for HalfCheetah. Since this is the only environment for which this was found to occur and we still see significant improvements in efficiency and stability, we apply LayerNorm to both the actor and the critic for all experiments in this paper (except where explicitly stated otherwise) towards our goal of improving training efficiency. We also expect the addition of LayerNorm to be universally introduced to off-policy reinforcement learning algorithms going forwards.
Appendix D Ablation of Critic Pre-Training
Figure 4: Ablation of critic pre-training (imitation only pre-training shown in red). This demonstrates the benefit of our proposed critic pre-training. Both implementations utilise LayerNorm. All lines show mean and standard deviation in normalised return at each timestep over 3 seeds.
We see that when pre-training the actor with behaviour cloning, the initial performance matches our proposed actor and critic pre-training, but quickly declines after pre-training due to the randomly initialised critic, matching the findings of Orsini et al., (2021). Therefore we see that the improved efficiency and stability from pre-training arises due to our combined actor and critic pre-training. This demonstrates that having a consistent actor and critic is essential for performance improvement.
Surprisingly, we also notice that in many cases the final performance is also more stable, even hundreds of thousands of updates after pre-training. As analysed in Fujimoto et al., (2022) and Chen and Jiang, (2019), for finite data regimes such as the offline setting, the Bellman equation can be satisfied by infinitely many suboptimal solutions. Additionally, $Q$-values trained by minimising the Bellman error are often empirically found be to be inaccurate (Schulman et al.,, 2018). We hypothesise that this additional stability could be occurring because the initial pre-training using the value error reduces the subset of possible solutions to those with lower value error when subsequently minimising the Bellman error on the finite offline dataset. However, since this benefit is auxiliary to our central focus of improving efficiency, we leave investigation of this effect to further work.
Appendix E Investigation into Empirical Bias Variance Tradeoff
Figure 5: Investigation into the empirical bias-variance tradeoff by varying $\lambda$ defined in equation 3 for TD3+BC on Hopper and HalfCheetah medium. All lines show mean and standard deviation in normalised return at each timestep over 3 seeds.
We investigate the bias-variance tradeoff described in section 3.2 in practice by empirically varying $\lambda$ for TD3+BC on the Hopper and HalfCheetah medium datasets. We find that while all values of $\lambda\in[0,1]$ provide efficiency benefits over no pre-training, the benefit is more significant for $\lambda=0$ (corresponding to the originally proposed value-error pre-training), likely because even for $\lambda=0.1$ the temporal difference component to the loss can have significant impact on the training dynamics, and the pre-training duration is not long enough for this bootstrapping loss to reach consistency. However, including some temporal difference component ($\lambda>0$) can help to smooth the transition with off-policy reinforcement learning as we see for HalfCheetah, while the inclusion of some Monte-Carlo component helps to improve the training efficiency.
We also see that stability on convergence for the Hopper environment is greater for smaller $\lambda$ and particularly apparent for $\lambda=0$, supporting our hypothesis in Section 4 and Appendix D that this additional stability follows from the use of the value error (rather than the temporal difference error) in pre-training.
Appendix F MuJoCo Medium-Replay and Full-Replay Dataset Experiments
We apply our pre-training approach proposed in Section 4 with an identical experimental implementation to that described in Section 5.1 to the medium-replay and full-replay datasets, shown below in Figures 6 and 7 respectively.
F.1 Medium-replay
The medium-replay dataset consists of 1M transitions from the replay buffer of an agent trained from random to medium performance. The influence of our pre-training approach is shown below.
We generally see similar efficiency and stability gains to those observed for the medium datasets in Figure 2. However, we see there are no efficiency gains to be made on the HalfCheetah environments given the baselines with our addition of LayerNorm converge so quickly. We also notice that the performance of Walker2d for EDAC is much cleaner, likely due to greater diversity of data helping to stabilise performance.
F.2 Full-replay
The full-replay dataset consists of 1M transitions from the replay buffer of an agent trained from random to expert performance. The influence of our pre-training approach is shown below.
Yet again we see similar efficiency and stability gains to those observed for the medium datasets in Figure 2 (with the exception of the HalfCheetah environments). In some cases, the efficiency gains may be further improved by optimisation of the pre-training duration and temporal difference component $\lambda$ from those that were originally selected for the medium datasets.
Appendix G Key Hyperparameters for Adroit Experiments
@@
The Adroit experiments were performed and evaluated as described in Section 5 of the main paper, with more detail provided below in Appendix H. Key hyperparameters for each algorithm are provided in Table 3 below. Full hyperparameter configurations (including those not provided below) are available in the config files of the provided CORL codebase (Tarasov et al.,, 2022). Where possible, hyperparameters were chosen to match previously published values for the Adroit environments, and otherwise their default implementation values. Crucially, where hyperparameters are shared between algorithms (such as between TD3+BC and TD3+BC+CQL) they were chosen to be equal, to investigate whether additional regularization on the critic/actor can improve over the exisiting tuned regularization on the actor/critic alone.
@@
To incorporate our additional regularization losses which may be of different scales to the existing losses, we utilise the normalisation strategy described in TD3+BC (Fujimoto and Gu,, 2021). Namely for primary loss $\alpha$ and additional auxiliary loss $\beta$ which may be of a different scale, we combine them as follows to more evenly balance the losses throughout training:
@@
where ||$\cdot$|| ⋅ | denotes the magnitude of the gradient-detached loss and $c$ is the regularization coefficient referred to as CQL/BC-regularizer provided in Table 3.
Finally, we note that for the behaviour cloning baseline and for the behaviour cloning regularization in both TD3+BC(+CQL) and EDAC+BC we use ‘hard’ behaviour cloning (using a mean-squared error objective). In the case of the BC baseline and EDAC+BC it would be possible to use ‘soft’ behaviour cloning as in Equation 6, but we found in both cases ‘hard’ behaviour cloning (using the sampled action from the Gaussian policy for EDAC) performed much better. However we still use ‘soft’ behaviour cloning for pre-training EDAC+BC to maintain the policy entropy in pre-training.
@@
Appendix H Evaluation Procedure for Adroit Experiements
@@
Evaluation for the Adroit experiments is carried out as described in Section 5. In particular, we train all algorithms using the standard hyperparameters provided above for 300k steps, and evaluate their performance by evaluating the agent performance every 10k steps between 200k and 300k steps. For the algorithms including pre-training, we pre-train for 200k steps to allow more than sufficient convergence of the pre-training stage (using $\lambda=0$). There are two motivations for this evaluation procedure. Firstly, we would like to measure the performance of the algorithms after a relatively short training time (i.e. reduced number of training steps) to provide a quantitative measure of the performance efficiency. Secondly, since the performance of the offline RL algorithms considered on this benchmark (and in general) are very unstable, online performance varies significantly between random and human performance during training. Therefore, to reduce the variance of the results and to incorporate performance stability, we do not take a single (or best) checkpoint, but rather average the performance of checkpoints taken every 10k steps between 200k and 300k training steps, each evaluated online for 10 episodes. This is more representative of real world performance off offline RL algorithms where the online performance may not be possible to evaluate during training.
@@
One additional consideration for evaluation resulted from the fact that we noticed that, aside from the pen environment, the human demonstrations in the human datasets were much longer than the truncation limit of the online truncation limit of the environments. The truncation limit for all environments (for the v1 environments other than pen) is set to 200 timesteps, while the maximum demonstration lengths are 300, 624 and 527 timesteps for the Door, Hammer and Relocate environments respectively. This partially explains the poor performance of previous algorithms on the non-pen environments in this benchmark (An et al.,, 2021), since the environment does not allow sufficient time to receive reward for successfully imitating the demonstrated behaviour, such as opening the door or hammering the nail, before truncating the epsiode! Therefore we adjust the truncation limit in these environments to match the maximum demonstration lengths. To compute the standard human normalized score (defined as $(\mathrm{agent\_score}-\mathrm{random\_score})/(\mathrm{human\_score}-\mathrm{random\_score})$, we maintain the same maximal human scores as provided in the d4rl benchmark, but adjust the minimal random scores by running the independent uniform random policy in the environments for the new truncation time limits. This gives rise to the new environment evaluation variables provided below in Table 4. Crucially, despite this improved evaluation procedure, we evaluate all algorithms considered using this procedure in an identical manner, to provide a fair comparison of algorithm performance that takes into account both efficiency and stability.
@@
pen $100$ $-162.09$ $3076.83$
door $300$ $-84.52$ $2880.57$
hammer $624$ $-856.83$ $12794.13$
relocate $527$ $-37.95$ $4233.88$
Appendix I Pre-Training Ablation on Adroit Environments
@@
door-human $0.6\pm 0.3$ $0.6\pm 0.6$ $0.9\pm 1.0$ $\mathbf{13.2\pm 3.8}$
@@
door-cloned $0.1\pm 0.4$ $0.0\pm 0.3$ $\mathbf{0.8\pm 0.7}$ $0.2\pm 0.5$
@@
relocate-cloned $0.0\pm 0.1$ $\mathbf{0.1\pm 0.0}$ $0.0\pm 0$ 0.0 $0.1\pm 0.1$
We see that both versions of each algorithm (with and without pre-training) perform similarly, although there appears to be a slight benefit from pre-training for EDAC+BC in some environments. This is likely partially just because there is little efficiency gain to be made from pre-training in this data-limited setting, and partly because the combined regularization is sufficient to keep the actor and critic consistent with each other and with the data distribution. Also both of the additional regularization components (behaviour cloning and CQL-style regularization) do not rely on temporal difference bootstrapping and therefore have similar efficiency to supervised learning and reduce the benefit of pre-training. However, a shorter pre-training period may demonstrate greater benefits from pre-training (since 200k steps is really more than required, providing additional time for non-pre-trained versions to learn), along with increased dataset size. Importantly, these hybrid algorithms, motivated by our pre-training approach, still demonstrate promising improvements in performance relative to the component algorithms in this data-limited setting, as demonstrated in Table 2.
We also note that even with our substantial efforts to make our evaluation procuedure as ‘fair’ as possible between algorithms (as described in Appendicies G and H) the variances (and therefore confidence intervals) of these Adroit results are still non-negligible due to the nature of the limited human data, the stochastic starting states of the environment, the high variance algorithms used, and our limited computation resources. However, our aim is not to show that any one algorithm is ‘best’, as this is dependent on a wide range of factors and is often entirely infeasible in general (Patterson et al.,, 2023). Indeed, the performance of these algorithms is generally comparable (as intuitively might be expected given the same limited behaviour data), and significantly improving on the behaviour policy is challenging. Rather, our results on this benchmark aim to demonstrate the idea that if the performance of an algorithm collapses after pre-training (or more generally, imitation learning gives rise to non-negligible performance but off-policy RL does not), this can be mitigated by introducing additional regularization towards the behaviour policy, and it is often more effective to regularize both the actor and the critic rather than just one of these components, as in Table 2.
Appendix J Performance Plots for Adroit Experiments

### v2

@@ G.2 Implementation Details
door-human $0.6\pm 0.7$ $6.6\pm 6.4$ $0.1\pm 0.1$ $0.2\pm 0.1$ $0.6\pm 0.6$ $\mathbf{13.2\pm 3.8}$
@@
door-cloned $0.3\pm 0.5$ $\mathbf{6.6\pm 3.6}$ $0.1\pm 0.5$ $1\pm 0.0 0.0$$0\pm 0.3$ $0.2\pm 0.5$
@@
Table 2: Normalized average returns ($\pm$ standard deviation), by averaging performance between 200k and 300k updates over 4 independent seeds. Our combined algorithms ensure that both the actor and critic are regularized to stay close to the behavior policy after pre-training, often giving greater performance than the component algorithms when learning from limited demonstrations.
G.3 Results and Analysis
@@
We find that behavior cloning (BC) provides a strong baseline, notably greatly outperforming previously quoted BC performances on these environments due to our addition of LayerNorm. We see that CQL and EDAC (with LayerNorm) perform reasonably with this evaluation approach using the same hyperparameters as specified in the original papers (kumar_conservative_2020; an_uncertainty-based_2021), although may benefit from a greater training budget. The original TD3+BC paper did not consider the Adroit environments, but we see comparable performances with relatively strong behavior cloning regularization (tuned with $\alpha=1$ for the pen environments and $\alpha=0.1$ for other environments). However, our additions of CQL regularization to TD3+BC and BC regularization to EDAC both generally lead to improved performance in these environments. In fact, we find that our new hybrid algorithms perform similarly using this evaluation procedure even without pre-training, due to their strong inherent regularization, as shown and discussed further in Appendix G.6. Performance plots in these environments for the human datasets are provided in Appendix G.7.
G.4 Key Hyperparameters for Adroit Experiments
@@
The Adroit experiments were performed and evaluated as described in Section 5 of the main paper, with more detail provided below in Appendix G.5. Key hyperparameters for each algorithm are provided in Table 3 below. Full hyperparameter configurations (including those not provided below) are available in the config files of the provided CORL codebase (tarasov2022corl). Where possible, hyperparameters were chosen to match previously published values for the Adroit environments, and otherwise their default implementation values. Crucially, where hyperparameters are shared between algorithms (such as between TD3+BC and TD3+BC+CQL) they were chosen to be equal, to investigate whether additional regularization on the critic/actor can improve over the exisiting tuned regularization on the actor/critic alone.
@@
To incorporate our additional regularization losses which may be of different scales to the existing losses, we utilize the normalization strategy described in TD3+BC (fujimoto_minimalist_2021). Namely for primary loss $\alpha$ and additional auxiliary loss $\beta$ which may be of a different scale, we combine them as follows to more evenly balance the losses throughout training:
@@
where $|\cdot|$ denotes the magnitude of the gradient-detached loss and $c$ is the regularization coefficient referred to as CQL/BC-regularizer provided in Table 3.
Finally, we note that for the behavior cloning baseline and for the behavior cloning regularization in both TD3+BC(+CQL) and EDAC+BC we use ‘hard’ behavior cloning (using a mean-squared error objective). In the case of the BC baseline and EDAC+BC it would be possible to use ‘soft’ behavior cloning as in Equation 6, but we found in both cases ‘hard’ behavior cloning (using the sampled action from the Gaussian policy for EDAC) performed much better. However we still use ‘soft’ behavior cloning for pre-training EDAC+BC to maintain the policy entropy in pre-training.
@@
G.5 Additional Discussion of Evaluation Procedure for Adroit Experiements
@@
As described in Section G.2, we train all algorithms using the standard hyperparameters provided for 300k steps, and evaluate their performance by evaluating the agent performance every 10k steps between 200k and 300k steps. For the algorithms including pre-training, we pre-train for 200k steps to allow more than sufficient convergence of the pre-training stage (using $\lambda=0$). There are two motivations for this evaluation procedure. Firstly, we would like to measure the performance of the algorithms after a relatively short training time (i.e. reduced number of training steps) to provide a quantitative measure of the performance efficiency. Secondly, since the performance of the offline RL algorithms considered on this benchmark (and in general) are very unstable, online performance varies significantly between random and human performance during training. Therefore, to reduce the variance of the results and to incorporate performance stability, we do not take a single (or best) checkpoint, but rather average the performance of checkpoints taken every 10k steps between 200k and 300k training steps, each evaluated online for 10 episodes. This is more representative of real world performance off offline RL algorithms where the online performance may not be possible to evaluate during training.
@@
One additional consideration for evaluation resulted from the fact that we noticed that, aside from the pen environment, the human demonstrations in the human datasets were much longer than the truncation limit of the online truncation limit of the environments. The truncation limit for all environments (for the v1 environments other than pen) is set to 200 timesteps, while the maximum demonstration lengths are 300, 624 and 527 timesteps for the Door, Hammer and Relocate environments respectively. This partially explains the poor performance of previous algorithms on the non-pen environments in this benchmark (an_uncertainty-based_2021), since the environment does not allow sufficient time to receive reward for successfully imitating the demonstrated behavior, such as opening the door or hammering the nail, before truncating the episode. Therefore we adjust the truncation limit in these environments to match the maximum demonstration lengths. To compute the standard human normalized score (defined as $(\mathrm{agent\_score}-\mathrm{random\_score})/(\mathrm{human\_score}-\mathrm{random\_score})$, we maintain the same maximal human scores as provided in the D4RL benchmark, but adjust the minimal random scores by running the independent uniform random policy in the environments for the new truncation time limits. This gives rise to the new environment evaluation variables provided below in Table 4. Crucially, despite this improved evaluation procedure, we evaluate all algorithms considered using this procedure in an identical manner, to provide a fair comparison of algorithm performance that takes into account both efficiency and stability.
@@
pen $100$ 162.09-162.09 $3076.83$
door $300$ 84.52-84.52 $2880.57$
hammer $624$ 856.83-856.83 $12794.13$
relocate $527$ 37.95-37.95 $4233.88$
G.6 Pre-Training Ablation on Adroit Environments
@@
door-human $\pm 0.3 0.6±0.6$$\pm 0.6$ $0.9\pm 1.0$ $\mathbf{13.2\pm 3.8}$
@@
door-cloned $1\pm 0.4 0.0$$0\pm 0.3$ $\mathbf{0.8\pm 0.7}$ $0.2\pm 0.5$
@@
relocate-cloned $0.0\pm 0.1$ $\mathbf{0.1\pm 0.0}$ $0.0\pm 0.0$ $0.1\pm 0.1$
We see that both versions of each algorithm (with and without pre-training) perform similarly, although there appears to be a slight benefit from pre-training for EDAC+BC in some environments. This is likely partially just because there is little efficiency gain to be made from pre-training in this data-limited setting, and partly because the combined regularization is sufficient to keep the actor and critic consistent with each other and with the data distribution. Also both of the additional regularization components (behavior cloning and CQL-style regularization) do not rely on temporal difference bootstrapping and therefore have similar efficiency to supervised learning and reduce the benefit of pre-training. However, a shorter pre-training period may demonstrate greater benefits from pre-training (since 200k steps is really more than required, providing additional time for non-pre-trained versions to learn), along with increased dataset size. Importantly, these hybrid algorithms, motivated by our pre-training approach, still demonstrate promising improvements in performance relative to the component algorithms in this data-limited setting, as demonstrated in Table 2.
We also note that even with our substantial efforts to make our evaluation procedure as ‘fair’ as possible between algorithms (as described in Appendices G.4 and G.5) the variances (and therefore confidence intervals) of these Adroit results are still non-negligible due to the nature of the limited human data, the stochastic starting states of the environment, the high variance algorithms used, and our limited computation resources. However, our aim is not to show that any one algorithm is ‘best’, as this is dependent on a wide range of factors and is often entirely infeasible in general (patterson_empirical_2023). Indeed, the performance of these algorithms is generally comparable (as intuitively might be expected given the same limited behavior data), and significantly improving on the behavior policy is challenging. Rather, our results on this benchmark aim to demonstrate the idea that if the performance of an algorithm collapses after pre-training (or more generally, imitation learning gives rise to non-negligible performance but off-policy RL does not), this can be mitigated by introducing additional regularization towards the behavior policy, and it is often more effective to regularize both the actor and the critic rather than just one of these components, as in Table 2.
G.7 Performance Plots for Adroit Experiments